Current spot exchange rate

$0.60 per £

Forward foreign exchange rate

$0.605 per £

Annualized interest rate on a 30-day dollar-denominated asset

15%

Annualized interest rate on a 30-day pound-denominated asset

7%

1. Does Covered Interest Parity hold?

2. If American investor invested in the U.K. What is the covered interest differential for the investor?
Answer #1

covered interest parity holds when

(1+id) = F(1+if)/S

where id is interest in domestic country and if is interest in foriegn country and F is forward rate, s is spot rate.

here F = $0.605 per £

S = $0.60 per £

so $ is domestic currency, USA is domestic country and id= 15%

and if = 7%

(1+id) = (1+0.15) = 1.15

F(1+if)/s = 0.605(1+0.07)/0.60 = 1.078

since (1+id) is not equal to F(1+if)/s, condition does not hold ture.

B) if american invests in UK,

borrow 1$ at 15%

purchase 1.66 pounds ( 0.6$ = 1 pound then 1$ = 1.66 pounds)

invest 1.66 pounds at 7% - get 1.78 pounds after 30 days

simultaneously sell 1.78 pounds at forward rate. get 1.07$ (1 pound = 0.605$)

loan repayment = 1+0.15=1.15$

return from investment = 1.07 - 1.15 = -0.08$.

-0.08$ is the interest differential